# A simpler and parallelizable $O(\sqrt{\log n})$-approximation algorithm for Sparsest Cut

### v4

@@ 4 Proof of Theorem 10
A $k$-hop extended matching cover is function ${\cal M}$ that maps vectors $u,u^{\prime}\in\mathbb{R}^{d}$ to a distribution over generalized matchings. It is skew-symmetric if sampling $M\sim{\cal M}(-u^{\prime},-u)$ and then reversing all paths in $M$ produces the same distribution as sampling $M\sim{\cal M}(u,u^{\prime})$. We define ${\tt size}_{\omega}({\cal M})=\tfrac{1}{n}\mathbb{E}_{M\sim{\cal M}({\cal N}_{\omega})}[|M|]$ where ${\cal M}({\cal N}_{\omega})$ denotes the distribution $(u,u^{\prime})\sim{\cal N}_{\omega},M\sim{\cal M}(u,u^{\prime})$. We say that ${\cal M}$ is $\sigma$-stretched (resp. $L$-long) if $\min\{\langle y-x,u\rangle,\langle y-x,u^{\prime}\rangle\}\geq\sigma$ (resp. $||y-x||^{2}\leq L$) for any $u,u^{\prime}\in\mathbb{R}^{d}$ and any nonviolating $(x,\ldots,y)\in{\tt supp}({\cal M}(u,u^{\prime}))$. We write $-~{\cal M}\subseteq\tilde{\cal M}$ for (coupled) extended matching covers $-~{\cal M},\tilde{\cal M}$ if $M\subseteq\tilde{M}$ for any $u,u^{\prime}$ and $M\sim{\cal M}(u,u^{\prime})$, $\tilde{M}\sim\tilde{\cal M}(u,u^{\prime})$.
@@
$\lambda_{1}$ $\delta_{1}$ (12a)
$\lambda_{2k}$ $\theta_{2k}\lambda_{k}^{2},\quad\theta_{2k}=\tfrac{1}{2}\left(\tfrac{1}{2}(\delta_{2k}-\delta_{k})\right)^{2/(1-\omega)}\quad\forall k\in\dot{\cal K}$ (12b)
@@
By plugging appropriate sequences $\{\delta_{k}\}$ and $\{\beta_{k}\}$ we can derive Theorem 10 as follows. (The proof of Lemma 17 is given in Appendix D).
@@
4.1 Proof of Theorem 16
Lemma 18.
There holds $\gamma_{x}\geq\lambda_{x}$.
@@ 4.1 Proof of Theorem 16
  • ()($\ast$)( ∗ )
Lemma 19.
There holds $\tau_{x}\geq\delta^{2/(1-\omega)}\lambda_{x}^{2}-2\varepsilon$ where $\varepsilon:=\exp\left(-\frac{\beta_{k}^{2}}{2(k+1)(1-\omega\rho)\Delta}\right)$.
Proof.
Define random variable $p_{1}$ and $p_{2}$ as follows:
We showed that
where we used (12b) and (13).
Appendix A Matrix multiplicative weights (MW) algorithm
It this section we review the method of Arora and Kale [AK16] for the checking the feasibility of system (${}^{\prime}$) consisting of constraints (${}^{\prime}$) and (2b)-(2e). The algorithm is given below.
1 for $t=1,2,\ldots,T$ do
2       compute
3       either output Fail or find “feedback matrix” $N^{(t)}$ of the form
$N^{(t)}={\tt diag}(y)+\sum_{p}f_{p}T_{p}+\sum_{S}z_{S}K_{S}-F$
where $f_{p}\geq 0$, $z_{S}\geq 0$, $\sum_{i}y_{i}+an^{2}\sum_{S}z_{S}\geq\alpha$, $F$ is a symmetric matrix with $F\preceq C$, and $N^{(t)}\bullet X^{(t)}\leq 0$.
Algorithm 5 MW algorithm.
It can be checked that $N^{(t)}\bullet X\geq 0$ for all feasible $X$ satisfying (${}^{\prime}$) (see [AK16]). Thus, matrix $N^{(t)}$ can be viewed as a cutting plane that certifies that $X^{(t)}$ is infeasible (or lies on the boundary of the feasible region).
The procedure at line 3 is called Oracle, and the maximum possible spectral norm $||N^{(t)}||$ of the feedback matrix is called the width of the oracle. This width will be denoted as $\rho$444 Note that [AK16] formulated the algorithm in terms of the “loss matrix” $M^{(t)}=-\frac{1}{\rho}N^{(t)}$ that satisfies $||M^{(t)}||\leq 1$. Namely, it used the update $W^{(t)}\!=\!\exp\left(-\bar{\eta}\sum_{r=1}^{t-1}M^{(r)}\right)$ with $\bar{\eta}\!=\!\rho\eta$, and set $\bar{\eta}\!=\!\frac{\varepsilon}{2\rho n}$ in their Theorems 4.4 and 4.6.
Theorem 20 ([AK16]).
A.1 Gram decomposition and matrix exponentiation
Theorem 21 ([AK16, Lemma 7.2]).
From now on we make the following assumption.
Assumption 1.
A.2 Oracle implementation
Proposition 22.
Suppose that $K_{S}\bullet\tilde{X}<\tfrac{an^{2}}{4}$. Then setting $y_{i}=-\tfrac{\alpha}{n}$ for all $i\in V$, $z_{S}=\tfrac{2\alpha}{an^{2}}$, $z_{S^{\prime}}=0$ for all $S^{\prime}\neq S$, and $F=0$ gives a valid output of the oracle with width $\rho=O(\tfrac{\alpha}{n})$ assuming that parameters $\tau,\gamma$ in Theorem 21 satisfy $\gamma\leq\tfrac{1}{2}$ and $\tau\leq\tfrac{a}{2}$.
Proof.
Denote $z_{ij}=||v_{i}-v_{j}||^{2}$ and $\tilde{z}_{ij}=||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$. We know that $\tilde{Z}:=\sum_{ij}\tilde{z}_{ij}=K_{S}\bullet\tilde{X}<\tfrac{an^{2}}{4}$ where the sum is over $i,j\in S$. Also, $|z_{ij}-\tilde{z}_{ij}|\leq\gamma(\tilde{z}_{ij}+\tau)$ for all $i,j$. This implies that
$K_{S}\bullet X=\sum_{ij}z_{ij}<\tilde{Z}+\gamma\tilde{Z}+\tfrac{n^{2}}{2}\gamma\tau\leq(1+\gamma)\tfrac{an^{2}}{4}+\tfrac{n^{2}}{2}\gamma\tau\leq\tfrac{an^{2}}{2}$
Note that $N^{(t)}=-\tfrac{\alpha}{n}I+\tfrac{2\alpha}{an^{2}}K_{S}$. We have $\sum_{i}y_{i}+an^{2}\sum_{S}z_{S}=n\cdot(-\tfrac{\alpha}{n})+an^{2}\cdot\tfrac{2\alpha}{an^{2}}=\alpha$ and $N^{(t)}\bullet X=-\tfrac{\alpha}{n}I\bullet X+\tfrac{2\alpha}{an^{2}}(K_{S}\bullet X)\leq-\tfrac{\alpha}{n}\cdot n+\tfrac{2\alpha}{an^{2}}\cdot\tfrac{an^{2}}{2}=0$, as desired. Also, $||N^{(t)}||\leq||-\tfrac{\alpha}{n}I||+||\tfrac{2\alpha}{an^{2}}K_{S}||=O(\tfrac{\alpha}{n})$. ∎
Now suppose that $\sum_{i,j\in S}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}=K_{S}\bullet\tilde{X}\geq\tfrac{an^{2}}{4}$. We will set $N^{(t)}_{ij}=0$ for all $(i,j)\notin S\times S$. When describing how to set the remaining entries of $N^{(t)}$, it will be convenient to treat $N^{(t)}$, $X$, $\tilde{X}$ etc. as matrices of size $|S|\times|S|$ rather than $n\times n$, and $y$ as a vector of size $|S|$. (Note that $X$ and $\tilde{X}$ are the submatrices of the original matrices, and ${\tt Tr}(X)\leq n$). With some abuse of terminology we will redefine $V=S$ and $n=|S|$, and refer to the original variables as $V_{\tt orig}$ and $n_{\tt orig}\in[n,2n]$. Thus, from now on we make the following assumption.
Assumption 2.
$||\tilde{v}_{i}||^{2}\leq 2$ for all $i\in V$ and $\sum_{i,j\in V:i<j}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}\geq\tfrac{an_{\tt orig}^{2}}{4}\geq\tfrac{an^{2}}{4}$.
Let us set $y_{i}=\frac{\alpha}{n}$ for all $i\in V$ and $z_{S}=0$ for all $S$, then $\sum_{i}y_{i}+an_{\tt orig}^{2}\sum_{S}z_{S}=\alpha$. Note that $N^{(t)}=\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F$ and $\tfrac{\alpha}{n}I\bullet X\leq\alpha$, so condition $(\sum_{p}f_{p}T_{p}-F)\bullet X\leq-\alpha$ would imply that $N^{(t)}\bullet X\leq 0$. Our goal thus becomes as follows.
We arrive at the specification of Oracle given in Section 2.1.
Appendix B Proof of Theorem 5
Here we repeat the argument in [${}^{+}$06]. For a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ denote $||f||_{p}=(\mathbb{E}_{u}[f^{p}(u)])^{1/p}$. Let $T_{\omega}f$ is the function $\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ defined via $(T_{\omega}f)(u)=\mathbb{E}_{\hat{u}\sim_{\omega}u}f(\hat{u})$. ($T_{\omega}$ is called the “Ornstein-Uhlenbeck operator”).
Theorem 23 ([Bor82]).
$||T_{\omega}f||_{q}\geq||f||_{p}$
$\geq||f||_{p}||T_{\omega}g||_{q}$ by reverse Hölder inequality
$\geq||f||_{p}||g||_{p}$ by Theorem 23
$=\delta^{2/(1-\omega)}$
Appendix C Proof of Lemma  6
(b)   Assume that condition (ii) is false. By a standard argument, Assumption 2 implies the following: there exist constants $c^{\prime}\in(0,c)$, $\beta>0$ and $\sigma>0$ such that with probability at least $\beta$ Algorithm 1 reaches line 9 and we have $w_{y}-w_{x}\geq\sigma$ for all $x\in A$, $y\in B$ (see [Kal07, Lemma 14]). Suppose that this event happens. We claim that in this case $|M|\geq\tfrac{1}{3}c^{\prime}n$. Indeed, suppose this is false. Let $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ be the sets of nodes involved in $M$ (with $|A^{\prime}|=|B^{\prime}|=|M|=k$). The total value of flow from $A$ to $B$ is at least $c^{\prime}\pi n$ (otherwise we would have terminated at line 5). The value of flow leaving $A^{\prime}$ is at most $|A^{\prime}|\cdot\pi\leq\tfrac{1}{3}c^{\prime}\pi n$. Similarly, the value of flow entering $B^{\prime}$ is at most $|B^{\prime}|\cdot\pi\leq\tfrac{1}{3}c^{\prime}\pi n$. Therefore, the value of flow from $A-A^{\prime}$ to $B-B^{\prime}$ is at least $c^{\prime}\pi n-2\cdot\tfrac{1}{3}c^{\prime}\pi n=\tfrac{1}{3}c^{\prime}\pi n$. For each edge $(x,y)\in M_{\tt all}$ with $x\in A-A^{\prime}$, $y\in B-B^{\prime}$ we have $||x-y||^{2}>\Delta$ (otherwise $M$ would not be a maximal matching in $M_{\tt short}$). Therefore,
But then the algorithm should have terminated at line 7 - a contradiction.
Appendix D Proof of Lemma 17
@@
Using (18), we conclude that if $\phi<\tfrac{1}{40}$ then $\pi_{K}\geq\Gamma_{K/2}(1-20\phi)\geq\tfrac{1}{8}\cdot\tfrac{1}{2}=\tfrac{1}{16}$.

### v5

@@ 4 Proof of Theorem 10
A $k$-hop extended matching cover is function ${\cal M}$ that maps vectors $u,u^{\prime}\in\mathbb{R}^{d}$ to a distribution over generalized matchings. It is skew-symmetric if sampling $M\sim{\cal M}(-u^{\prime},-u)$ and then reversing all paths in $M$ produces the same distribution as sampling $M\sim{\cal M}(u,u^{\prime})$. We define ${\tt size}_{\omega}({\cal M})=\tfrac{1}{n}\mathbb{E}_{M\sim{\cal M}({\cal N}_{\omega})}[|M|]$ where ${\cal M}({\cal N}_{\omega})$ denotes the distribution $(u,u^{\prime})\sim{\cal N}_{\omega},M\sim{\cal M}(u,u^{\prime})$. We say that ${\cal M}$ is $\sigma$-stretched (resp. $L$-long) if $\min\{\langle y-x,u\rangle,\langle y-x,u^{\prime}\rangle\}\geq\sigma$ (resp. $||y-x||^{2}\leq L$) for any $u,u^{\prime}\in\mathbb{R}^{d}$ and any nonviolating $(x,\ldots,y)\in{\tt supp}({\cal M}(u,u^{\prime}))$. We write ${\cal M}\subseteq\tilde{\cal M}$ for (coupled) extended matching covers ${\cal M},\tilde{\cal M}$ if $M\subseteq\tilde{M}$ for any $u,u^{\prime}$ and $M\sim{\cal M}(u,u^{\prime})$, $\tilde{M}\sim\tilde{\cal M}(u,u^{\prime})$.
@@
[…]
@@
By plugging appropriate sequences $\{\delta_{k}\}$ and $\{\beta_{k}\}$ we can derive Theorem 10 as follows.
@@
Using (17), we conclude that if $\phi<\tfrac{1}{40}$ then $\pi_{K}\geq\Gamma_{K/2}(1-20\phi)\geq\tfrac{1}{8}\cdot\tfrac{1}{2}=\tfrac{1}{16}$. ∎
4.1 Proof of Theorem 16
Lemma 18.
$\gamma_{x}\geq\lambda_{x}$.
Proof.
  • ()($\ast$)( ∗ )
Lemma 19.
$\tau_{x}\geq\delta^{2/(1-\omega)}\lambda_{x}^{2}-2\varepsilon$ where $\varepsilon:=\exp\left(-\frac{\beta_{k}^{2}}{2(k+1)(1-\omega\rho)\Delta}\right)$.
Proof.
Define random variable $p_{1}$ and $p_{2}$ as follows:
We showed that
where we used (12b) and (13).
Appendix A Matrix multiplicative weights (MW) algorithm
It this section we review the method of Arora and Kale [AK16] for the checking the feasibility of system (2) consisting of constraints (2a) and (2b)-(2e). The algorithm is given below.
1 for $t=1,2,\ldots,T$ do
2       compute
3       either output Fail or find “feedback matrix” $N^{(t)}$ of the form
$N^{(t)}={\tt diag}(y)+\sum_{p}f_{p}T_{p}+\sum_{S}z_{S}K_{S}-F$
where $f_{p}\geq 0$, $z_{S}\geq 0$, $\sum_{i}y_{i}+\xi n^{2}\sum_{S}z_{S}\geq\alpha$, $F$ is a symmetric matrix with $F\preceq C$, and $N^{(t)}\bullet X^{(t)}\leq 0$.
4
Algorithm 5 MW algorithm.
It can be checked that $N^{(t)}\bullet X\geq 0$ for all feasible $X$ satisfying (2) (see [AK16]). Thus, matrix $N^{(t)}$ can be viewed as a cutting plane that certifies that $X^{(t)}$ is infeasible (or lies on the boundary of the feasible region).
The procedure at line 3 is called Oracle, and the maximum possible spectral norm $||N^{(t)}||$ of the feedback matrix is called the width of the oracle. This width will be denoted as $\rho$555 Note that [AK16] formulated the algorithm in terms of the “loss matrix” $M^{(t)}=-\frac{1}{\rho}N^{(t)}$ that satisfies $||M^{(t)}||\leq 1$. Namely, it used the update $W^{(t)}\!=\!\exp\left(-\bar{\eta}\sum_{r=1}^{t-1}M^{(r)}\right)$ with $\bar{\eta}\!=\!\rho\eta$, and set $\bar{\eta}\!=\!\frac{\varepsilon}{2\rho n}$ in their Theorems 4.4 and 4.6.
Theorem 20 ([AK16, Theorems 4.4 and 4.6]).
A.1 Gram decomposition and matrix exponentiation
Theorem 21 ([AK16, Lemma 7.2]).
From now on we make the following assumption.
Assumption 1.
A.2 Oracle implementation
Proposition 22.
Suppose that $K_{S}\bullet\tilde{X}<\tfrac{\xi n^{2}}{4}$. Then setting $y_{i}=-\tfrac{\alpha}{n}$ for all $i\in V$, $z_{S}=\tfrac{2\alpha}{\xi n^{2}}$, $z_{S^{\prime}}=0$ for all $S^{\prime}\neq S$, and $F=0$ gives a valid output of the oracle with width $\rho=O(\tfrac{\alpha}{n})$ assuming that parameters $\tau,\gamma$ in Theorem 21 satisfy $\gamma\leq\tfrac{1}{2}$ and $\tau\leq\tfrac{\xi}{2}$.
Proof.
Denote $z_{ij}=||v_{i}-v_{j}||^{2}$ and $\tilde{z}_{ij}=||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$. We know that $\tilde{Z}:=\sum_{ij}\tilde{z}_{ij}=K_{S}\bullet\tilde{X}<\tfrac{\xi n^{2}}{4}$ where the sum is over $i,j\in S$. Also, $|z_{ij}-\tilde{z}_{ij}|\leq\gamma(\tilde{z}_{ij}+\tau)$ for all $i,j$. This implies that
$K_{S}\bullet X=\sum_{ij}z_{ij}<\tilde{Z}+\gamma\tilde{Z}+\tfrac{n^{2}}{2}\gamma\tau\leq(1+\gamma)\tfrac{\xi n^{2}}{4}+\tfrac{n^{2}}{2}\gamma\tau\leq\tfrac{\xi n^{2}}{2}$
Note that $N^{(t)}=-\tfrac{\alpha}{n}I+\tfrac{2\alpha}{\xi n^{2}}K_{S}$. We have $\sum_{i}y_{i}+\xi n^{2}\sum_{S}z_{S}=n\cdot(-\tfrac{\alpha}{n})+\xi n^{2}\cdot\tfrac{2\alpha}{\xi n^{2}}=\alpha$ and $N^{(t)}\bullet X=-\tfrac{\alpha}{n}I\bullet X+\tfrac{2\alpha}{\xi n^{2}}(K_{S}\bullet X)\leq-\tfrac{\alpha}{n}\cdot n+\tfrac{2\alpha}{\xi n^{2}}\cdot\tfrac{\xi n^{2}}{2}=0$, as desired. Also, $||N^{(t)}||\leq||-\tfrac{\alpha}{n}I||+||\tfrac{2\alpha}{\xi n^{2}}K_{S}||=O(\tfrac{\alpha}{n})$. ∎
Now suppose that $\sum_{i,j\in S}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}=K_{S}\bullet\tilde{X}\geq\tfrac{\xi n^{2}}{4}$. We will set $N^{(t)}_{ij}=0$ for all $(i,j)\notin S\times S$. When describing how to set the remaining entries of $N^{(t)}$, it will be convenient to treat $N^{(t)}$, $X$, $\tilde{X}$ etc. as matrices of size $|S|\times|S|$ rather than $n\times n$, and $y$ as a vector of size $|S|$. (Note that $X$ and $\tilde{X}$ are the submatrices of the original matrices, and ${\tt Tr}(X)\leq n$). With some abuse of terminology we will redefine $V=S$ and $n=|S|$, and refer to the original variables as $V_{\tt orig}$ and $n_{\tt orig}\in[n,2n]$. Thus, from now on we make the following assumption.
Assumption 2.
$||\tilde{v}_{i}||^{2}\leq 2$ for all $i\in V$ and $\sum_{i,j\in V:i<j}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}\geq\tfrac{\xi n_{\tt orig}^{2}}{4}\geq\tfrac{\xi n^{2}}{4}$.
Let us set $y_{i}=\frac{\alpha}{n}$ for all $i\in V$ and $z_{S}=0$ for all $S$, then $\sum_{i}y_{i}+\xi n_{\tt orig}^{2}\sum_{S}z_{S}=\alpha$. Note that $N^{(t)}=\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F$ and $\tfrac{\alpha}{n}I\bullet X\leq\alpha$, so condition $(\sum_{p}f_{p}T_{p}-F)\bullet X\leq-\alpha$ would imply that $N^{(t)}\bullet X\leq 0$. Our goal thus becomes as follows.
We arrive at the specification of Oracle given in Section 2.1.
Appendix B Proof of Theorem 5
Here we repeat the argument in [MOR+06]. For a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ denote $||f||_{p}=(\mathbb{E}_{u}[f^{p}(u)])^{1/p}$. Let $T_{\omega}f$ is the function $\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ defined via $(T_{\omega}f)(u)=\mathbb{E}_{\hat{u}\sim_{\omega}u}f(\hat{u})$. ($T_{\omega}$ is called the “Ornstein-Uhlenbeck operator”).
Theorem 23 ([Bor82]).
$||T_{\omega}f||_{q}\geq||f||_{p}$
$\geq||f||_{p}||T_{\omega}g||_{q}$ by reverse Hölder inequality
$\geq||f||_{p}||g||_{p}$ by Theorem 23
$=\delta^{2/(1-\omega)}$